• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Emily

    Log entry time 08:04:08 on August 12, 2006

    Entry number 179416

    keyword=Owl Shift/Run Summary

    Shift Summary

    Continuing Kinematics Point 18 production runs. 0:00 - 8uA, LD2, 5%, 3/4" and 3" C doors in 0:50 - Left spectrometer angle was off by ~3 degrees, and affects runs back to 2566. 3:05 - 15uA,requested. Doubled prescale values. 3:25 - 12uA, requested because radiator temp too high 3:30 - 8uA, requested because radiator temp not dropping enough. 4:35 - Closed JT valves partway because reaching 25g/s limit. There seems to be no response. Refer to log 179370, M.O.'s description. 7:05 - Compton High Voltage Alarm

    Run Summary

    Run # # Events Current

    2577/29035 1.9M, .7M, 8uA, LD2, 5%, 3/4" and 3" C doors in.. 2578/20936 1.9M, .7M 2579/20937 1.9M, .7M 2580/20938 1.9M, .7M 2581/20939 1.9M, .7M 2582/20940 1.9M, .7M 2583/20941 1.9M, .7M 2584/20942 1.9M, .7M 2585/20943 1.9M, .7M 2586/20944 1.9M, .7M 2587/20945 1.9M, .7M 2588/20946 1.9M, .7M 2589/20947 1.9M, .7M

    2590/20948 1.9uA, 1.3uA, 15uA, LD2, 5%, 3/4" and 3" C doors in. 2591/20949 1.5M, 1.0M

    2592/20950 1.9M, .7M, 8uA, LD2, 5%, 3/4" and 3" C doors in. 2593/20951 1.9M, .7M 2594/20952 1.9M, .7M 2595/20953 1.9M, .7M 2596/20954 1.9M, .7M 2597/20955 1.9M, .7M 2598/20956 1.9M, .7M 2599/20957 1.9M, .7M 2600/20958 1.9M, .7M 2601/20959 1.9M, .7M 2602/20960 1.9M, .7M 2603/20961 1.9M, .7M 2604/20961 1.9M, .7M 2605/20963 1.9M, .7M 2606/20964 1.9M, .7M 2607/20965 1.9M, .7M 2608/20966 1.9M, .7M